# OODFace: Benchmarking Robustness of Face Recognition under Common Corruptions and Appearance Variations

### v1

@@ 5.3 Commercial API Evaluation Results
We evaluate the robustness of three commercial FR services (Aliyun, iFLYTEK, and Tencent) on LFW-C and LFW-V. The underlying mechanisms and training data of these APIs are entirely unknown to us. We follow their original threshold ranges and determine the optimal threshold for each dataset. The results are shown in Tab. 4 and Tab. 5 .
Common Corruptions Evaluation. For LFW-C, our first observation is that although the designed corruptions appear mild and have minimal impact on human recognition, they significantly impair the performance of commercial FR systems, often returning “No face detected.” Based on this, we compute the Rejection Rate (RR), representing the proportion of rejected samples. We find that nearly all categories of corruption have cases where the algorithm fails, with the most significant impact in the Data & Processing category, especially for Salt-and-Pepper Noise, where Aliyun’s API completely fails and gains 100.00% RR. We also compute Accepted Samples Accuracy (ASA) and Actual Accuracy (AA), representing the success rate among accepted samples and the overall samples, respectively. While Aliyun’s ASA is relatively high, this is due to its high RR, resulting in a low proportion of accepted samples. Ultimately, Tencent’s API achieves the best RR, highest ASA, and AA, but there is still a noticeable decline in accuracy compared to clean data. Given the substantial detection rejection rates, we conclude that current FR systems need significant improvements in robustness against corruption.
Models Accuracy
Variations Aliyun iFLYTEK Tencent
None (clean) 99.65 97.99 99.75
Age Age- 96.60 93.47 96.45
Age+ 96.60 93.50 96.59
Facial Expression Mouth-close 96.67 93.80 96.47
Mouth-open 96.55 93.68 96.55
Eye-close 96.71 93.23 96.64
Eye-open 96.67 94.10 96.72
Rotation Rotation-left 96.67 94.23 96.71
Rotation-right 96.71 94.10 96.67
Accessories Bangs&Glasses 98.06 93.91 98.85
Makeup 99.13 96.60 99.28
Average 97.27 94.42 97.33
Table 5: Accuracy of 3 commercial APIs on LFW-V.
Appearance Variations Evaluation. For LFW-V, no detection rejection occurs as seen in LFW-C, so we report the direct accuracy results. Appearance variations also cause performance degradation, but the impact is smaller compared to corruptions. Among the APIs, Tencent achieves the highest average accuracy, while Aliyun performs better in the Age and Facial Expression. The robustness of these services remains highly correlated with their clean performance, aligning with the conclusions in Sec. 5.2.
Insight 5: Faces with several corruption types fail to be detected by APIs, causing FR systems to malfunction in several OOD scenarios, highlighting the vulnerability of commercial FR systems.
@@ 5.4 Extended Experiments on Face Masks
In the field of FR, face masks typically refer to physical disguises used to deceive or obscure the identity of the wearer, preventing the system from accurately recognizing the individual [55]. Masks can replicate the appearance of a real face, thereby bypassing the system’s security measures. As an extension, we conduct physical experiments on masks, creating five different types made from various materials, as shown in Fig. 7. In the experiments, we first test whether the masks could be correctly identified, and then further explore the impact of corruptions and variations on recognition. Our findings indicate that masks are also affected by OOD scenarios, exhibiting a different category of performance degradation compared to real faces. We provide additional experimental data and further analysis in Appendix E, along with details of the mask fabrication process.
Insight 6: Face masks are also affected by OOD scenarios, exhibiting a different performance degradation trend compared to real faces.
5.5 Extended Experiments on VLMs
Figure 8: Testing GPT-4o mini for FR. GPT-4o mini demonstrates robust FR capabilities. (More results in Appendix G.)
@@ 5.5 Extended Experiments on VLMs
To explore better solutions for OOD scenarios in FR, we investigate the use of Vision-Language Models (VLMs) for FR, a previously unexplored approach. As shown in Fig. 8, we find that GPT-4o-mini [1] demonstrates robust FR capabilities and can identify specific types of corruption present in the images, suggesting that VLMs could be a solution to the challenge of OOD data.
Insight 7: VLMs can compare facial similarities even in OOD scenarios, suggesting that leveraging the generalization capabilities of large models could be a promising solution for handling OOD cases.
5.6 Extended Experiments on Potential Defenses
We further explore potential defenses for improving robustness in Appendix F. Following [52], we test 10 robust models using input transformation such as R&P [50], Bit-Red [51], and adversarial training such as PGD-AT [32], TRADES [53]. However, the improvements are limited.
Insight 8: Existing defense strategies fail to address OOD scenarios, and enhancing FR robustness remains an open challenge for future research.
6 Conclusion
@@
In this paper, we introduce OODFace, a comprehensive benchmark for OOD robustness in FR, systematically designed with 20 common corruptions across 5 categories and 10 appearance variations across 4 categories. By augmenting public datasets, we establish three robustness benchmarks: LFW-C/V, CFP-FP-C/V, and YTF-C/V. We conduct extensive evaluations on 19 FR models and 3 commercial APIs, along with additional experiments towards face masks, VLMs, and defense strategies. We hope that our comprehensive benchmarks, detailed analysis, and insightful findings will aid in understanding the robustness of FR models against OOD scenarios, and provide valuable guidance for future improvements in FR model robustness.
@@ A.1 Implementation Details of Corruptions
Figure A.1: Display of face masks and robotic arm for data collection
@@
A.2. Appearance Variations Implementation Details

### v2

@@ 5.3 Commercial API Evaluation Results
We evaluate three commercial FR services (Aliyun, iFLYTEK, and Tencent) on LFW-C/V. We follow their original threshold ranges and determine the optimal threshold for each dataset. The results of LFW-C are shown in Tab. 4. We leave the result of LFW-V in Appendix Tab. B.1.
Common Corruptions Evaluation. For LFW-C, the corruption data significantly impair the performance of commercial FR systems, often returning “No face detected.” Based on this, we compute the Rejection Rate (RR), representing the proportion of rejected samples. We find that nearly all categories of corruption have cases where the algorithm fails, especially for Salt-and-Pepper Noise, where Aliyun’s API completely fails and gains 100.00% RR. We also compute Accepted Samples Accuracy (ASA) and Actual Accuracy (AA), representing the success rate among accepted samples and the overall samples, respectively. While Aliyun’s ASA is relatively high, this is due to its high RR, resulting in a low proportion of accepted samples. Ultimately, Tencent’s API achieves the best RR, highest ASA, and AA, but there is still a noticeable decline in accuracy.
Insight 5: Commercial FR APIs exhibit catastrophic failures under corruptions. Excessive Rejection Rates paradoxically inflate Accepted Sample Accuracy but cause complete system breakdowns in identity verification, highlighting the vulnerability of commercial FR systems.
@@ 5.4 Extended Experiments on Face Masks
Face masks typically refer to physical disguises used to obscure the identity of the wearer, preventing the system from accurately recognizing the individual [59]. As an extension, we conduct physical experiments on masks, creating five types made from various materials, as shown in Fig. 7 (Detailed result and analysis in Appendix G).
Insight 6: Face masks exhibit material-dependent vulnerability patterns under OOD scenarios, with degradation trends differing from real faces. This suggests a potential approach for spoof detection.
6 Experiments Result on Potential Solutions
In Sec. 5, FR models suffer performance degradation on the OODs. In this section, we explore potential solutions from two perspectives: applying potential defenses in Sec. 6.1; leveraging VLMs to address OOD challenges in Sec. 6.2.
6.1 Experiments on Defense Methods
Figure 8: Display of restoration methods as potential defenses.
Following [56], we test 10 robust models using input transformation such as R&P [54], Bit-Red [55], and adversarial training such as PGD-AT [35], TRADES [57]. As Tab. A.1 in Appendix E, the improvements are limited. Further, following [12], we explored three restoration methods [51, 63, 31] respectively based on GANs, Transformers, and Diffusion Models as potential defenses, and report the result in Appendix E, Tab. A.2. Experiments show that restoration improves performance for certain OOD categories (e.g., Data & Processing), especially for models with weaker robustness, highlighting its potential. However, restoration can also interfere with feature extraction, leading to performance degradation in some OOD categories. Detailed analyses are provided in Appendix E.
Open-source Commercial
Models LLaVA-NeXT InternVL2.5 Qwen-VL-Plus GPT4o-mini
Accuracy 52.04 50.00 87.76 98.98
Table 5: VLMs FR performance on corruption data.
6.2 Experiments on VLMs
@@
To further explore solutions for OOD scenarios, we investigate the use of Vision-Language Models for FR, a previously unexplored approach. We find that GPT-4o-mini [1] demonstrates robust FR capabilities and can identify specific types of corruption in the images (see Fig. D.1 in Appendix. F). We further test more open-source (LLaVA-NeXT-8B, InternVL2.5-8B) and commercial (Qwen-VL-Plus) VLMs on the corruption data, with results shown in Tab. 5, and additional analysis in Appendix. F.
Insight 7: Existing defenses fail to fully mitigate OOD scenarios. VLMs demonstrate robust FR potential under OODs, suggesting a solution for robust FR. However, only commercial models achieve strong performance, the architectures and training processes remain opaque, posing challenges for deployment. Enhancing FR robustness remains an open challenge for future research.
7 Discussion and Conclusion
@@
In this paper, we introduce OODFace, a comprehensive benchmark for OOD robustness in FR, systematically designed with 20 common corruptions across 5 categories and 10 appearance variations across 4 categories. By augmenting public datasets, we establish three robustness benchmarks: LFW-C/V, CFP-FP-C/V, and YTF-C/V. We conduct extensive evaluations on 19 FR models and 3 commercial APIs, along with additional experiments towards face masks, VLMs, and defense strategies. Experimental results demonstrate that FR models suffer severe performance degradation under OOD scenarios, while existing strategies fail to fully mitigate these challenges. Based on the results and our insights, we outline potential research directions:
  • 1)
    Adaptive Defense Orchestration: Future defenses could focus on adaptive defense orchestration, enabling OOD-aware defense selection.
  • 2)
    Decoupled Feature Restoration: Developing decoupled feature restoration modules could help prevent feature distortions introduced by defense strategies.
  • 3)
    Robust FR Models: Training with our comprehensive OOD dataset could be a promising approach to improve generalization and robustness of FR models.
  • 4)
    VLMs for FR: It is essential to enhance the interpretability of the robust FR performance before exploring the design of specialized FR-VLMs or integrating VLMs into FR pipelines to enhance robustness. Above all, privacy concerns must be carefully addressed.
We hope that our comprehensive benchmarks, detailed analysis and insights will aid in understanding the robustness of FR models against OOD scenarios, and provide guidance for future improvements in FR model robustness.
@@ A.1 Implementation Details of Corruptions
Figure A.1: t-SNE visualization of feature distributions before and after applying different OOD simulations to the models.
Restoration Models Open-Source Model Eval Model Architecture Eval Optimization Loss Eval Average
GFPGAN L & W Brightness -0.68 -0.53 -0.72 -0.47 -0.40 -0.67 -0.20 -0.08 -0.30 -0.35 0.02 -0.42 -0.50 -0.43 -0.42 -0.38 -0.38 -0.35 -0.37 -0.40
Contrast -2.22 -0.67 -3.18 -1.18 -0.92 -1.43 -0.65 -0.92 -1.75 -1.77 -1.57 -1.43 -1.38 -1.80 -2.13 -2.17 -1.40 -1.67 -1.82 -1.58
Saturate -0.70 -0.60 -1.07 -0.58 -0.30 -0.67 -0.25 -0.02 -0.23 -0.13 -0.23 -0.25 -0.45 -0.32 -0.35 -0.57 -0.47 -0.23 -0.37 -0.41
Fog -4.32 0.92 -1.33 -0.32 -0.63 -2.53 -0.67 0.30 0.05 2.68 2.25 1.13 1.32 0.17 0.10 -0.12 0.58 -0.57 -0.22 -0.06
Snow 2.63 2.70 -0.80 -1.33 -1.35 -3.00 -0.82 0.07 0.90 1.57 1.62 1.32 0.65 0.62 -0.05 -0.30 -1.20 -0.42 -0.70 0.11
Sensor Defocus Blur -6.60 10.15 3.47 -0.62 -6.82 -10.63 -5.43 -1.63 -0.15 1.77 2.13 0.28 1.93 -3.03 -0.48 -2.42 0.05 -3.42 -4.07 -1.34
Color Shift -0.25 0.08 -0.53 -0.12 -0.07 -0.18 -0.13 0.05 0.02 -0.08 -0.08 -0.13 0.02 -0.15 -0.17 -0.18 -0.02 -0.05 -0.03 -0.11
Pixelate -0.90 0.03 -0.30 -0.22 -0.22 -0.63 -0.25 0.03 0.03 -0.20 0.23 0.00 -0.17 -0.28 -0.35 -0.62 -0.15 -0.30 -0.25 -0.24
Movement Motion Blur -2.90 4.57 1.85 -0.20 -1.57 -3.30 -1.13 -0.35 0.12 1.02 1.05 0.08 0.88 -0.72 -0.62 -1.32 0.20 -0.98 -1.17 -0.24
Zoom Blur -0.33 -0.95 -0.82 -0.48 -0.58 -1.28 -0.53 -0.08 -0.33 -0.33 -0.17 -0.25 -0.30 -0.52 -0.40 -0.57 -0.47 -0.52 -0.62 -0.50
Facial Distortion -2.32 13.18 6.07 0.48 1.10 -4.32 0.73 2.98 1.62 4.15 3.88 1.62 2.00 2.47 1.97 1.60 0.97 -0.03 0.78 2.05
D & P Gaussian Noise 2.27 23.77 10.10 1.42 5.12 -0.53 5.22 12.15 7.10 21.33 18.88 17.20 9.05 7.88 7.80 9.27 3.48 2.95 5.08 8.92
Impulse Noise␣️ 1.20 21.55 7.38 0.63 0.98 -1.12 2.68 6.60 4.73 14.52 13.62 14.83 8.95 7.15 4.77 3.88 2.03 0.98 2.25 6.19
Shot Noise 4.10 25.27 14.48 1.67 5.13 -0.40 6.10 12.15 9.73 25.68 23.03 22.98 13.33 11.28 12.18 14.78 5.45 4.30 7.77 11.53
Speckle Noise 2.80 22.55 10.65 1.02 2.07 -0.82 3.37 6.35 7.18 19.85 18.67 17.68 11.33 7.48 8.33 9.22 4.28 3.03 5.37 8.44
Salt Pepper Noise 11.12 23.15 25.63 4.93 9.75 -2.20 14.57 30.03 17.43 27.10 25.07 36.52 30.33 24.47 34.47 33.28 9.63 16.37 21.08 20.67
Jpeg Compression -0.90 0.42 -0.40 -0.13 -0.22 -0.50 -0.22 -0.02 -0.27 -0.05 -0.07 -0.10 -0.02 -0.17 -0.25 -0.43 -0.35 -0.15 -0.20 -0.21
Occlusion Random Occlusion -5.28 0.42 -0.30 -1.92 -1.52 -2.13 -1.83 -0.05 -0.98 -1.15 -1.28 -0.50 -1.18 -1.00 -1.02 -0.63 -1.45 -0.82 -1.00 -1.24
Frost 1.75 8.23 1.50 0.48 0.80 -2.68 1.02 3.40 0.32 4.12 3.80 1.20 1.72 2.03 1.15 2.42 0.28 0.47 0.37 1.70
Spatter -0.83 1.17 -0.85 -1.17 -1.23 -2.02 -0.75 -0.47 -0.95 0.02 -0.15 -1.15 -0.17 -0.33 -0.87 -1.23 -1.30 -0.95 -1.08 -0.75
Average -0.12 7.77 3.54 0.09 0.46 -2.05 1.04 3.52 2.21 5.99 5.54 5.53 3.87 2.74 3.18 3.18 0.99 0.88 1.54 2.63
CodeFormer L & W Brightness -0.75 -0.47 -0.57 -0.35 -0.38 -0.62 -0.12 0.00 -0.38 -0.28 -0.20 -0.42 -0.35 -0.25 -0.22 -0.42 -0.45 -0.23 -0.28 -0.35
Contrast -2.97 -1.17 -3.55 -1.58 -1.55 -2.25 -1.07 -1.22 -2.33 -2.87 -2.93 -2.58 -2.00 -2.55 -2.32 -2.98 -2.17 -2.15 -2.33 -2.24
Saturate -0.55 -0.63 -0.77 -0.18 -0.07 -0.53 -0.18 0.12 -0.20 -0.08 -0.22 -0.30 -0.33 -0.08 -0.28 -0.30 -0.35 -0.08 -0.08 -0.27
Fog -4.82 0.22 -1.07 -1.03 -1.97 -3.93 -1.57 -0.98 -0.15 2.03 1.90 0.20 0.68 -0.27 0.32 -0.67 0.27 -1.27 -1.07 -0.69
Snow 2.38 0.67 -1.35 -0.83 -0.72 -2.13 -0.20 0.55 0.45 0.65 0.38 0.63 -0.40 0.08 0.07 -0.67 -0.62 -0.53 -0.40 -0.10
Sensor Defocus Blur -10.43 9.23 2.95 -2.73 -8.28 -12.57 -5.92 -3.63 -1.48 1.97 0.92 -0.50 1.47 -3.78 -2.88 -5.92 0.08 -5.17 -6.23 -2.79
Color Shift -0.58 -0.53 -0.67 -0.27 -0.07 -0.37 -0.17 -0.03 -0.13 -0.35 -0.27 -0.33 -0.13 -0.22 -0.33 -0.23 -0.25 -0.12 -0.18 -0.28
Pixelate -1.17 -0.20 -0.75 -0.33 -0.25 -0.70 -0.30 -0.02 -0.18 -0.38 -0.28 -0.08 -0.30 -0.13 -0.55 -0.65 -0.17 -0.20 -0.33 -0.37
Movement Motion Blur -2.60 5.30 2.18 0.12 -0.77 -2.23 -0.40 0.33 0.48 1.80 1.55 0.55 1.45 0.10 0.00 -0.70 0.75 -0.52 -1.03 0.34
Zoom Blur -1.10 -1.20 -0.68 -0.60 -0.38 -1.13 -0.38 0.08 -0.28 -0.37 -0.47 -0.30 -0.23 -0.27 -0.27 -0.50 -0.27 -0.28 -0.37 -0.47
Facial Distortion -4.20 11.10 4.80 -0.07 0.60 -4.45 0.20 2.33 0.87 3.80 2.72 1.27 1.73 2.03 1.73 1.15 0.87 -0.53 -0.17 1.36
D & P Gaussian Noise 1.72 23.10 9.92 0.70 3.65 -2.00 4.73 11.03 6.20 20.50 18.12 16.82 8.87 7.32 7.00 8.50 3.13 2.17 4.02 8.18
Impulse Noise␣️ 0.20 20.30 6.60 0.25 0.30 -1.82 2.45 6.00 3.88 13.63 12.37 14.35 8.48 6.62 3.77 2.95 1.22 0.48 1.55 5.45
Shot Noise 3.00 22.60 12.87 1.28 4.20 -1.35 6.08 11.32 8.77 24.75 21.55 22.53 12.95 11.00 11.32 14.02 4.92 3.97 6.90 10.67
Speckle Noise 0.70 14.73 8.00 0.55 1.72 -1.32 3.63 6.08 6.17 17.57 15.60 15.42 9.18 6.83 7.05 8.52 3.05 2.38 4.53 6.86
Salt Pepper Noise 13.02 22.82 27.52 4.90 10.20 0.23 16.62 28.92 18.63 30.47 27.28 38.97 31.73 27.35 37.52 37.30 10.53 16.85 22.22 22.27
Jpeg Compression -1.42 -0.35 -0.80 -0.47 -0.72 -1.20 -0.47 -0.48 -0.50 -0.55 -0.53 -0.37 -0.15 -0.58 -0.83 -0.88 -0.80 -0.55 -0.62 -0.65
Occlusion Random Occlusion -4.90 -0.48 -1.10 -2.30 -2.23 -2.52 -1.97 -1.07 -1.98 -1.87 -2.02 -1.43 -2.50 -2.05 -2.23 -1.55 -2.22 -1.78 -2.30 -2.03
Frost 0.42 6.23 0.68 -0.15 -0.17 -3.78 0.98 2.83 -0.23 3.00 2.58 0.92 1.05 1.02 0.47 1.93 -0.08 -0.02 -0.63 0.90
Spatter -1.12 -0.12 -1.30 -0.93 -0.57 -1.38 -0.50 0.08 -0.95 -0.58 -0.85 -1.42 -0.78 -0.48 -0.68 -0.92 -1.03 -0.65 -0.63 -0.78
Average -0.76 6.56 3.15 -0.20 0.13 -2.30 1.07 3.11 1.83 5.64 4.86 5.20 3.52 2.58 2.93 2.90 0.82 0.59 1.13 2.25
DiffBIR L & W Brightness -0.62 -0.85 -0.65 -0.85 -0.72 -1.35 -0.63 -0.37 -0.85 -0.75 -0.53 -0.57 -0.67 -0.65 -0.70 -0.85 -1.08 -0.68 -0.75 -0.74
Contrast -8.72 -8.15 -11.62 -8.48 -10.35 -11.98 -8.75 -9.60 -11.42 -13.25 -12.02 -11.77 -10.32 -12.00 -9.23 -10.18 -10.98 -10.37 -10.92 -10.53
Saturate -0.53 -0.68 -1.17 -0.38 -0.22 -0.67 -0.30 0.03 -0.22 -0.30 -0.60 -0.30 -0.47 -0.32 -0.42 -0.35 -0.47 -0.17 -0.25 -0.41
Fog -8.93 -3.57 -6.55 -7.10 -9.38 -12.87 -9.40 -7.70 -7.25 -5.78 -4.93 -5.85 -5.58 -7.52 -4.77 -5.97 -5.80 -8.10 -7.72 -7.09
Snow 2.98 2.12 -0.30 -0.90 -0.37 -2.30 -0.70 0.57 0.70 1.25 0.52 0.73 0.28 0.77 0.38 0.27 -0.40 -0.28 -0.37 0.26
Sensor Defocus Blur -12.47 5.60 -2.38 -6.88 -14.02 -18.75 -11.93 -8.30 -5.57 -3.82 -4.63 -4.97 -3.20 -9.10 -7.28 -10.23 -4.75 -10.52 -11.47 -7.61
Color Shift -0.75 -0.62 -0.72 -0.47 -0.38 -0.68 -0.32 -0.25 -0.53 -0.53 -0.83 -0.63 -0.35 -0.42 -0.55 -0.73 -0.58 -0.35 -0.33 -0.53
Pixelate -3.28 -2.50 -3.12 -2.25 -3.17 -4.45 -2.58 -2.37 -2.83 -3.22 -2.30 -2.28 -2.45 -2.57 -2.37 -3.08 -2.53 -2.75 -3.03 -2.80
Movement Motion Blur -3.53 2.58 -0.33 -2.42 -5.47 -7.88 -3.98 -2.90 -2.77 -1.33 -1.47 -2.83 -1.52 -3.37 -2.23 -3.02 -1.78 -3.97 -3.97 -2.75
Zoom Blur -1.40 -1.93 -1.85 -1.53 -1.75 -2.88 -1.57 -0.85 -1.50 -1.78 -1.47 -1.20 -1.43 -1.57 -1.08 -2.05 -1.35 -1.53 -1.42 -1.59
Facial Distortion -3.77 11.52 5.28 -1.65 -2.48 -8.35 -2.17 0.10 -0.43 2.00 1.73 -0.32 0.88 0.72 0.50 -0.90 -0.27 -2.72 -2.20 -0.13
D & P Gaussian Noise -2.23 17.75 5.15 -2.63 -0.43 -7.07 0.37 7.43 2.52 15.75 12.98 12.80 5.10 3.25 3.45 4.53 -0.67 -1.58 0.05 4.03
Impulse Noise␣️ -0.02 17.28 4.92 -0.12 0.17 -2.28 1.68 5.80 3.28 12.28 11.65 13.30 7.23 6.03 3.87 3.28 0.78 0.27 1.78 4.80
Shot Noise -1.57 16.57 7.48 -3.22 -1.53 -7.47 1.05 6.57 2.87 17.75 15.75 16.12 7.17 4.40 5.93 8.48 0.07 -1.92 1.57 5.06
Speckle Noise -2.10 15.47 5.30 -2.30 -2.82 -6.08 -0.58 2.27 2.52 13.90 12.82 12.62 6.08 2.73 3.88 5.05 0.32 -1.45 1.28 3.63
Salt Pepper Noise 12.10 14.23 21.53 2.53 7.05 -4.92 11.70 28.65 15.30 23.10 20.87 33.62 26.52 22.95 32.68 32.18 8.00 14.00 19.33 17.97
Jpeg Compression -2.08 -1.10 -1.67 -1.28 -1.67 -2.05 -1.17 -1.15 -1.62 -1.62 -1.38 -1.60 -1.45 -1.33 -1.53 -1.92 -1.43 -1.55 -1.58 -1.54
Occlusion Random Occlusion -4.28 0.60 -0.80 -2.28 -2.45 -3.58 -2.58 -1.20 -1.90 -1.98 -1.88 -1.52 -2.08 -1.63 -1.55 -1.37 -2.38 -2.13 -2.12 -1.95
Frost 1.98 8.08 1.27 -0.37 -0.45 -4.52 0.67 2.10 0.08 4.02 3.43 1.18 0.67 1.77 0.95 2.63 -0.20 0.05 -0.35 1.21
Spatter -0.32 1.53 -0.30 -0.83 -0.30 -1.23 -0.32 0.32 -0.67 0.38 -0.13 -0.47 0.05 0.20 -0.23 -0.28 -0.77 -0.27 -0.32 -0.21
Average -1.98 4.70 0.97 -2.17 -2.54 -5.57 -1.58 0.96 -0.51 2.80 2.38 2.80 1.22 0.12 0.98 0.78 -1.31 -1.80 -1.14 -0.05
Table A.2: Accuracy differences after implementing three defense methods on LFW-C. Green squares indicate improvement, while red squares indicate a decrease.
A.2 Implementation Details of Variations